Can someone help me on this?
I need to write a proportion that is equivalent to the statement: If 3 gallons of gasoline cost $5.37, then 8 gallons will cost $14.32.
Thanks,
Lyn

Answer by fractalier(6550)

A proportion is
a is to b as c is to d
or in fraction form
a/b = c/d
so we have
3/5.37 = 8/14.32
or equivalently
3/8 = 5.37/14.32